a. Calculate the fraction of methyl isonitrile molecules that have an energy of 160.0 kJ or greater at 504 K .

answer:

f1 =

2.61×10?17

b. Calculate this fraction for a temperature of 519 K .

c. What is the ratio of the fraction at 519 K to that at 504 K ?
Answer #1

b) According to Boltzmann distribution

The fraction of moles (f) = e-Ea/RT

Where Ea is the activation energy, R is the universal gas constant and T is the absolute temperature.

i.e. ln(f2/f1) = Ea/R * (1/T1 - 1/T2)

Here, f1 = 2.61*1017 molecules = 2.61*1017/6.023*1023 mol = 4.33*10-7 mol

i.e. ln(f2/4.33*10-7 mol) = 160*103 J/(8.314 J mol-1 K-1) * (1/504 - 1/519) K

i.e. ln(f2/4.33*10-7 mol) = 1.1036

i.e. f2/4.33*10-7 mol = e1.1036 = 3.015

i.e. f2 = 4.33*10-7 * 3.015 * 6.023*1023 molecules = 7.86*1017 molecules

(c) The ratio of the fraction at 519 K to 504 K

= f2/f1 = 3.015